1. Find "x" for these 2 problems
a)9^2x=2*(9^x)+3
b)3*(3^x)+9*(3^-x)=28
ive graduated 2years ago and i dont remember how to do those 2 problems
id like to know how to solve them step by step thanks
yours
brian

Answer by scott8148(5880)

a) let z=9^x ... z^2=2z+3 ... z^2-2z-3=0 ... (z-3)(z+1)=0 ... z=3, z=-1
... the negative value is not realistic ... 3=9^x ... x=1/2

b) multiplying by 3^x gives 3*(3^x)^2+9=28(3^x) ... the method of solution is similar to a)